OpenStudy (anonymous):

Can anyone help me, i don't seem to understand. let f(x)=2x g(x)=x^2+2 and h(x)=-4x+3 how do you find the composite function for. (Fog) and (Gof)

OpenStudy (anonymous):

I'll provide multiple choice answers provided if needed.

OpenStudy (anonymous):

\[f(x)=2x\]\[g(x)=x^2+2\]\[f(g[x]) = 2(x^2+2)\]

OpenStudy (anonymous):

because you "put" g(x) inside f(x)

OpenStudy (anonymous):

So it's 4x^2+2?

OpenStudy (usukidoll):

distribute the 2 2(x^2+2) 2(x^2)+(2)(2)

OpenStudy (anonymous):

So this question is actually two separate questions one to find the (fog) and (gof) but have the same multiple choices which are: A. 2x^2+4 B. 4x^2 +2 C. -8x+6 D. -4x^2-5 My guesses for (Gof) is B (Fog) is A Is this correct?

OpenStudy (usukidoll):

F o g is A we have 2x^2+4 I gotta do G o F one sec

OpenStudy (usukidoll):

(2x)^2+2 4x^2+2 is B

OpenStudy (usukidoll):

yeah it's right

OpenStudy (anonymous):

Ok just making sure, thank you so much!